In a photoelectric experiment set up, photons of energy 5 eV falls on the cathode having work function 3 eV. (a) If the saturation current is i = 4μA for intensity 10-5 W / m2 , then plot a graph between anode potential and current. (b) Also draw a graph for intensity of incident radiation 2 x 10-5 W/m2

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Hello Student,
Please find the answer to your question
a = Initial Uranium atom
(a – x ) = Uranium toms left
(a – x ) = a (1 / 2)n
and n = t / t1/2 = 1.5 x 109 / 4.5 x 109 = 13
∴ a – x = a (1/2)1/3
⇒ a / a-x = 1 / ( 1 / 2)1/3 = 21/3 / 1 = 1. 26
⇒ x / a – x = 1.26 – 1 = 0 . 26
Thanks
Navjot Kalra
askIITians Faculty